OpenStudy (anonymous):

Help ? Choose the equation of the line passing through the point (-4, -2) and parallel to y = 1/4x + 2 a. y = 4x - 1 b. y = 4x - 3 c. 1/4x - 3 d. 1/4x - 1

OpenStudy (emilylauren):

alright :) so if it's a parallel line the slope will remain the same, right?

OpenStudy (emilylauren):

The slope of the first line is 1/4. This means the slope to our new line is also going to be 1/4.

OpenStudy (emilylauren):

does this equation look familiar? y-(point Y)= m(x-(point X)) m=slope which is 1/4 point y=-2 point x=-4 just plug in the numbers to find the answer :)

OpenStudy (emilylauren):

y-(-2)=(1/4)(x-(-4))